# Rethinking Misalignment in Vision-Language Model Adaptation from a Causal Perspective

### v1

@@ 1 Introduction
Figure 1: The motivating examples of the two-level misalignment. (a) The cosine similarities of the image and some texts in the embedding space of CLIP. (b) Accuracy of base classes and new classes with varying training epochs on the DTD dataset.
However, not all categories in downstream tasks have annotated data available for prompt tuning. As a result, the distribution of training classes (i.e., base classes) often differs from that of testing classes (i.e., new classes), which constitutes the second misalignment, i.e., data misalignment. Researchers generally agree that prompt tuning potentially leads to the overfitting of the CLIP model to the base classes, deviating CLIP from its original behavior and thus hindering its generalization to new classes [7, 8]. To explore whether the overfitting holds, we conduct prompt tuning based on MaPLe [6] on the DTD dataset [9]. We record the performance of the base classes and the new classes under different training epochs. As shown in Figure 1(b), as the number of training epochs increases, the performance on base classes gradually improves; however, the performance on the new classes first increases and then decreases. The phenomenon is consistent with the definition of overfitting.
@@ 3.1 Problem Formulation
CLIP [1] is an impressive vision-language model comprised of an image encoder and a text encoder, capable of embedding images and text into a shared multi-modal latent space. Through contrastive learning, paired image-text embeddings are optimized for maximum cosine similarity, while unpaired ones are minimized. The intuition behind such design is based on a fundamental assumption: any paired images and texts are expected to form a tight semantic correspondence, as they contain similar content. By pre-training on a large-scale dataset consisting of 400 million image-text pairs, CLIP [1] demonstrates strong generalization in zero-shot and few-shot downstream tasks.
@@
$p(y=c|x)=\frac{exp(sim(v,w_{c})/\tau)}{\Sigma_{c^{\prime}=1}^{C}exp(sim(v,w_{c^{\prime}})/\tau)}$ (1)
@@ 3.2 Problem Analysis: Causal Perspective
Firstly, we define each variable in Figure 2. $D$ represents the data used in the pre-training stage. Generative factors, the fundamental semantic elements that control the generation of images and texts, can determine the content of both images and texts. In a specific task, these generative factors can be divided into two subsets: the set of task-relevant factors $G_{r}$ and the set of task-irrelevant factors $G_{i}$. Taking the pet classification task as an example, task-relevant generative factors determine features such as coat colors, body shapes, and ear types, which are highly relevant to identifying pet breeds. Conversely, task-irrelevant generative factors influence semantics which do not provide helpful information for pet classification. For a downstream task, $X$ represents the image space, while $Y$ corresponds to the true labels of these images. $\hat{Y}$ represents the predicted labels obtained through the learning. Additionally, $S$ denotes the semantics within the images $X$ closely relevant to the current task.
@@
By training on diverse data, CLIP acquires relatively comprehensive visual and language knowledge. Therefore, it can be assumed that through the learning of $D$, we obtain comprehensive generative factors that constitute the entire visual space. Since the learning process of pre-training is not specific to any particular downstream task, these generative factors include not only task-related but also task-irrelevant ones, i.e., $D\to G_{i}$ and $D\to G_{r}$, for a specific task. By combining the generative factors, any image $x\in X$ can be generated, regardless of whether it has appeared in the pre-training dataset $D$. Images typically include not only content related to their category under a particular task but usually other content as well, hence $G_{r}\to X$ and $G_{i}\to X$. However, only task-related generative factors can determine the true label $Y$ of images, i.e., $G_{r}\to Y$.
@@
Given that the training data $D$ for CLIP is invariant, in Figure 2(a), the path $X\leftarrow G_{i}\leftarrow D\to G_{r}\to Y$ is blocked, resulting in $X$ being associated with $Y$ solely through $G_{r}$, i.e., $X\leftarrow G_{r}\to Y$. When performing downstream classification tasks, we aim to model the relationship between $X$ and $\hat{Y}$ that arises from $G_{r}$. As shown in Figure 2(b), $X\to\hat{Y}$ represents the objective of our modeling. $G_{i}$ and $G_{r}$ are unobservable and easily confused in practice. Therefore, we perform soft prompt tuning to help capture the information of $G_{r}$ and eliminate the influence of $G_{i}$. However, the estimated task-related factors are not always accurate. Furthermore, due to the data misalignment, the task-relevant generative factors estimated based on the training data of base classes are not necessarily the task-relevant factors for new classes. $\hat{G_{i}}$ denotes the set of task-irrelevant generative factors that are incorrectly retained, and the set of factors that are task-relevant for the base classes but task-irrelevant for the new classes. Since we incorrectly assume that $\hat{G_{i}}$ is task-relevant, $\hat{G_{i}}$ influences $\hat{Y}$, i.e., $\hat{G_{i}}\to\hat{Y}$. When the learned prompt template is overfitted to the base classes, generative factors in $\hat{G}$ will further increase, hindering the estimation of the true causal relationship between $X$ and $\hat{Y}$.
@@
Figure 3: The framework of our proposed CDC. $t^{m}$ denotes a single template, while $p_{1},p_{2},...,p_{d}$ represent tokens in the template. Different colors indicate diverse templates. The text encoder and the image encoder are frozen, and only the tokens in the prompt templates are learnable.
@@ 4 Methodology
Guided by the front-door criterion, we propose CDC, comprising two essential components: VSD and DSTC. Specifically, VSD focuses on decoupling the semantics represented by diverse templates. DSTC aims to generate the classification scores via Equatnion 2 based on the decoupled semantics. We will elaborate on the details of these two components in Section 4.1 and Section 4.2, respectively.
@@ 4.1 Visual-Language Dual Semantic Decoupling
Intuitively, the embedding vector $w_{c}^{m}$ from the $m$-$th$ template can be regarded as a sample to be classified. The objective of Equation 3 is to maximize the entropy of the classification results, resulting in a uniform distribution. The classification results are obtained using the embedding of all other templates. $\mathcal{L}_{de}$ aims to enhance the semantic distinctions among different templates, thereby encouraging each template to maintain a unique semantic.
@@
where $w^{0}$ denotes the embeddings of the hand-crafted template $t^{0}$ obtained via the original CLIP model. See Appendix B for the detailed templates.
Via VSD, we obtain distinct semantic sets that contain decoupled semantics. Different classes have varying values on the same semantic set, so we consider the values of $s$ in Equation 2 to be the text embeddings of each class on all semantic sets, i.e, $s\in\{w^{m}_{c}\}_{m=1,c=1}^{M,C}$.
@@ 4.2 Decoupled Semantic Trusted Classification
To implement Equation 2, we need to clarify how to estimate the $P(\hat{Y}|do(x))$ based on the decoupled semantics. Firstly, the probability $P(s|x)$ that a sample $x$ has semantic $s$ is obtained by calculating the cosine similarity between the representation $v$ of $x$ and $s$, i.e., $P(s|x)=N(sim(s,v))$, where $N$ denotes the normalization operation.
After determining the possible values of $s$, we clarify how to estimate the second term in Equation 2. Let $x^{\prime}$ represent the samples from the training set and $s=w_{c}^{m}$, then: (i) if $c\neq y$, $P(y|x^{\prime},s)=0$. By updating the template, we expect each class $c$ to carry unique semantics $w_{c}^{m}$. Therefore, it cannot be assumed that the specific semantics of $c$ is attributable to another class; (ii) if $c=y$ but $x^{\prime}$ does not belong to the $c$-$th$ class, $P(y|x^{\prime},s)=0$; (iii) if $c=y$ and $x^{\prime}$ belong to the $c$-$th$ class, we maximize the similarity between $s$ and $x^{\prime}$ by prompt tuning so that the semantics of $s$ and $x^{\prime}$ are as consistent as possible. As $x^{\prime}$ contains both task-relevant and task-irrelevant semantics, and its task-relevant semantic is decoupled to $w_{c}^{1},w_{c}^{2},...,w_{c}^{M}$ via VSD, $s$ is the subset of $x^{\prime}$. Therefore, the second term in Equation 2 shrinks to $P(y|s)$.
Ideally, if $s$ contains and only contains the category-specific semantic information that explicitly supports the classification of sample $x^{\prime}$ into $c$, $P(y=c|s)=1$, and $P(y\neq c|x^{\prime},s)=0$. However, because of the data misalignment, while applying the learned templates to the new classes, this ideal scenario does not happen. In reality, the qualities of the learned semantics vary for different templates and classes, allowing the value of $P(y|s)$ to alter, thereby resulting in significant differences in the uncertainty of classification results derived from different templates. To solve this issue, we propose to estimate the uncertainty [29, 30] of each prediction from diverse templates via Dempster-Shafer theory [31, 32]. Specifically, we consider $e_{c}^{m}=h(sim(w_{c}^{m},v))$ as the evidence of classifying the feature $v$ as belonging to the $c$-$th$ class under the $m$-$th$ semantic set, where $h$ represents a function that maps the cosine similarity to a non-negative value. We assume that there exists a Dirichlet distribution $\mathbf{\alpha}^{m}=\left[\alpha_{1}^{m},\alpha_{2}^{m},...\alpha_{C}^{m}\right]$, where $\alpha_{c}^{m}=e_{c}^{m}+1$. This formulation allows us to quantify the belief that the feature $v$ belongs to class $c$ as well as the uncertainty associated with the prediction as:
@@
By sequentially fusing the results of all semantic sets, we can obtain the final classification results. Let $\mathcal{F}$ denote the fusion operation of Equation 7, then Equation 2 can be formally expressed as:
$P({\hat{Y}=c}|do({x}))=\mathcal{F}_{m=1}^{M-1}(b_{c}^{m},u^{m},b_{c}^{m+1},u^{m+1}).$ (8)
Accordingly, during the training stage, the cross-entropy loss which is used for learning the prompts is adjusted to the trusted cross-entropy loss to model the uncertainty information. For each training sample, the trusted cross-entropy can be formalized as:
@@ 4.3 Overall Architecture
where $\beta$ and $\gamma$ denote the hyper-parameters to tune the influence of $\mathcal{L}_{de}$ and $\mathcal{L}_{con}$, respectively. Refer to Section A for the detailed training and testing pipeline.
@@ 5 Experiments
Following previous works [5, 6], we conduct experiments to evaluate our proposed method with three different settings. These settings encompass the base-to-new setting as well as two out-of-distribution (OOD) settings, i.e., the cross-dataset setting and the cross-domain setting. Refer to Appendix C for a detailed overview of the evaluation protocol.
@@ 5.1 Experimental Settings
Experimental Details. We follow the experimental settings of the baseline method MaPLe. Specifically, we utilize a pre-trained CLIP with ViT-B/16 as the visual encoder. The number of learnable tokens is fixed at 2, whereas the prompt depth varies, being 9 for the base-to-new setting and 3 for the OOD setting. The learning rate is 0.035, and the batch size is 4. All models are trained using an SGD optimizer on an NVIDIA 3090 GPU. Our proposed CDC introduces three additional hyperparameters: $\beta$ and $\gamma$, which represent the weights for $\mathcal{L}_{de}$ and $\mathcal{L}_{con}$, respectively, and $M$, which denotes the number of prompts. Furthermore, different augmentation methods have a notable impact on the model’s performance. We set $\beta=5$, $\gamma=0.01$, and $M=4$ in the base-to-new setting, and $\beta=3$, $\gamma=0.01$, and $M=8$ in the OOD setting. For a detailed analysis of the influence of varying hyper-parameters and augmentation methods, please refer to Appendix D. We report the accuracies of base classes and novel classes, along with their harmonic mean (HM), averaged over 3 runs.

### v2

@@ 1 Introduction
Figure 1: (a) A motivating example of task misalignment, illustrating the cosine similarities between an image and various text descriptions in the embedding space of CLIP. (b) A motivating experiment on data misalignment, showing the accuracy trends for base and new classes across different training epochs on the DTD dataset.
However, not all downstream tasks have available annotated data for prompt tuning. As a result, inconsistency exists between the training and testing data, constituting the second misalignment, i.e., data misalignment. The inconsistency in data misalignment can arise due to two main reasons: (1) Label Inconsistency: The training and testing classes do not completely overlap. For instance, some classes present in the training data might not appear in the testing data and vice versa. We refer to the classes that appear in both training and testing as base classes and the classes that appear only in testing as new classes. (2) Distribution Inconsistency: Even if they share the same class names, the distributions of the classes in the training and testing data may differ, resulting in distribution inconsistency. In such cases, the testing classes are essentially also new classes. Researchers generally agree that prompt tuning potentially leads to the overfitting of the CLIP model to the base classes, deviating CLIP from its original behavior and thus hindering its generalization to new classes [7, 8]. To explore whether the overfitting holds, we conduct prompt tuning based on MaPLe [6] on the DTD dataset [9]. We record the performance of the base classes and the new classes under different training epochs. As shown in Figure 1(b), as the number of training epochs increases, the performance on base classes gradually improves; however, the performance on the new classes first increases and then decreases. The phenomenon is consistent with the definition of overfitting.
@@ 3.1 Problem Formulation
CLIP [1] is an impressive vision-language model comprised of an image encoder and a text encoder, capable of embedding images and texts into a shared multi-modal latent space. Through contrastive learning, paired image-text embeddings are optimized for maximum cosine similarity, while unpaired ones are minimized. The intuition behind such design is based on a fundamental assumption: any paired images and texts are expected to form a tight semantic correspondence, as they contain similar content. By pre-training on a large-scale dataset consisting of 400 million image-text pairs, CLIP [1] demonstrates strong generalization in zero-shot and few-shot downstream tasks.
@@
$p(y=c|x)=\frac{exp(sim(v,w_{c})/\tau)}{\Sigma_{c^{\prime}=1}^{C}exp(sim(v,w_{c^{\prime}})/\tau)},$ (1)
@@ 3.2 Problem Analysis: Causal Perspective
Firstly, we define each variable in Figure 2. $\mathcal{D}$ represents the data used in the pre-training stage. Generative factors, the fundamental semantic elements that control the generation of images and texts, can determine the content of both images and texts. In a specific task, these generative factors can be divided into two subsets: the set of task-relevant factors $G_{r}$ and the set of task-irrelevant factors $G_{i}$. Taking the pet classification task as an example, task-relevant generative factors determine features such as coat colors, body shapes, and ear types, which are highly relevant to identifying pet breeds. Conversely, task-irrelevant generative factors influence semantics which do not provide helpful information for pet classification. For a downstream task, $X$ represents the image space, while $Y$ corresponds to the true labels of these images. $\hat{Y}$ represents the predicted labels obtained through the learning. Additionally, $S$ denotes the semantics within the images $X$ closely relevant to the current task.
@@
By training on diverse data, CLIP acquires relatively comprehensive visual and language knowledge. Therefore, it can be assumed that through the learning of $\mathcal{D}$, we obtain comprehensive generative factors that constitute the entire visual space. Since the learning process of pre-training is not specific to any particular downstream task, these generative factors include not only task-related but also task-irrelevant ones, i.e., $\mathcal{D}\to G_{i}$ and $\mathcal{D}\to G_{r}$, for a specific task. By combining the generative factors, any image $x\in X$ can be generated, regardless of whether it has appeared in the pre-training dataset $\mathcal{D}$. Images typically include not only content related to their category under a particular task but usually other content as well, hence $G_{r}\to X$ and $G_{i}\to X$. However, only task-related generative factors can determine the true label $Y$ of images, i.e., $G_{r}\to Y$.
@@
Given that the training data $\mathcal{D}$ for CLIP is invariant, in Figure 2(a), the path $X\leftarrow G_{i}\leftarrow\mathcal{D}\to G_{r}\to Y$ is blocked, resulting in $X$ being associated with $Y$ solely through $G_{r}$, i.e., $X\leftarrow G_{r}\to Y$. When performing downstream classification tasks, we aim to model the relationship between $X$ and $\hat{Y}$ that arises from $G_{r}$. As shown in Figure 2(b), $X\to\hat{Y}$ represents the objective of our modeling. $G_{i}$ and $G_{r}$ are unobservable and easily confused in practice. Therefore, we perform soft prompt tuning to help capture the information of $G_{r}$ and eliminate the influence of $G_{i}$. However, the estimated task-related factors are not always accurate. Furthermore, due to the data misalignment, the task-relevant generative factors estimated based on the training data of base classes are not necessarily the task-relevant factors for new classes. $\hat{G_{i}}$ denotes the set of task-irrelevant generative factors that are incorrectly retained, and the set of factors that are task-relevant for the base classes but task-irrelevant for the new classes. Since we incorrectly assume that $\hat{G_{i}}$ is task-relevant, $\hat{G_{i}}$ influences $\hat{Y}$, i.e., $\hat{G_{i}}\to\hat{Y}$. When the learned prompt template is overfitted to the base classes, generative factors in $\hat{G}$ will further increase, hindering the estimation of the true causal relationship between $X$ and $\hat{Y}$.
@@
Figure 3: Framework of CDC. $t^{m}$ denotes a single template, while $p_{1},p_{2},...,p_{d}$ represent tokens in the template. Different colors indicate diverse templates. “fuse” refers to the process of generating the final classification results from multiple template results as shown in Equation (8). The text encoder and the image encoder are frozen, and only the tokens in the prompt templates are learnable.
@@ 4 Methodology
Guided by the front-door criterion, we propose CDC, comprising two essential components: VSD and DSTC. Specifically, VSD focuses on decoupling the semantics represented by diverse templates. DSTC aims to generate the classification scores via Equation (2) based on the decoupled semantics. We will elaborate on the details of these two components in Section 4.1 and Section 4.2, respectively.
@@ 4.1 Visual-Language Dual Semantic Decoupling
Intuitively, the embedding vector $w_{c}^{m}$ from the $m$-$th$ template can be regarded as a sample to be classified. The objective of Equation (3) is to maximize the entropy of the classification results, resulting in a uniform distribution. The classification results are obtained using the embedding of all other templates. $\mathcal{L}_{de}$ aims to enhance the semantic distinctions among different templates, thereby encouraging each template to maintain a unique semantic.
@@
where $w^{0}$ denotes the embeddings of the hand-crafted template $t^{0}$ obtained via the original CLIP model. See Appendix A for the detailed templates.
Via VSD, we obtain distinct semantic sets that contain decoupled semantics. Different classes have varying values on the same semantic set, so we consider the values of $s$ in Equation (2) to be the text embeddings of each class on all semantic sets, i.e, $s\in\{w^{m}_{c}\}_{m=1,c=1}^{M,C}$.
@@ 4.2 Decoupled Semantic Trusted Classification
To implement Equation (2), we need to clarify how to estimate the $P(\hat{Y}|do(x))$ based on the decoupled semantics. Firstly, the probability $P(s|x)$ that a sample $x$ has semantic $s$ is obtained by calculating the cosine similarity between the representation $v$ of $x$ and $s$, i.e., $P(s|x)=N(sim(s,v))$, where $N$ denotes the normalization operation.
After determining the possible values of $s$, we clarify how to estimate the second term in Equation (2). Let $x^{\prime}$ represent the samples from the training set and $s=w_{c}^{m}$, then: (i) if $c\neq y$, $P(y|x^{\prime},s)=0$. By updating the template, we expect each class $c$ to carry unique semantics $w_{c}^{m}$. Therefore, it cannot be assumed that the specific semantics of $c$ is attributable to another class; (ii) if $c=y$ but $x^{\prime}$ does not belong to the $c$-$th$ class, $P(y|x^{\prime},s)=0$; (iii) if $c=y$ and $x^{\prime}$ belong to the $c$-$th$ class, we maximize the similarity between $s$ and $x^{\prime}$ by prompt tuning so that the semantics of $s$ and $x^{\prime}$ are as consistent as possible. As $x^{\prime}$ contains both task-relevant and task-irrelevant semantics, and its task-relevant semantic is decoupled to $w_{c}^{1},w_{c}^{2},...,w_{c}^{M}$ via VSD, $s$ is the subset of $x^{\prime}$. Therefore, the second term in Equation (2) shrinks to $P(y|s)$.
Ideally, if $s$ contains and only contains the category-specific semantic information that explicitly supports the classification of sample $x^{\prime}$ into $c$, $P(y=c|s)=1$, and $P(y\neq c|x^{\prime},s)=0$. However, because of the data misalignment, while applying the learned templates to the new classes, this ideal scenario does not happen. In reality, the qualities of the learned semantics vary for different templates and classes, allowing the value of $P(y|s)$ to alter, thereby resulting in significant differences in the uncertainty of classification results derived from different templates. To solve this issue, we propose to estimate the uncertainty [29, 30] of each prediction from diverse templates via Dempster-Shafer theory [31, 32]. Specifically, we consider $e_{c}^{m}=h(sim(w_{c}^{m},v))$ as the evidence of classifying the feature $v$ as belonging to the $c$-$th$ class under the $m$-$th$ semantic set, where $h$ represents a function that maps the cosine similarity to a non-negative value. We assume that there exists a Dirichlet distribution $\mathbf{\alpha}^{m}=\left[\alpha_{1}^{m},\alpha_{2}^{m},...,\alpha_{C}^{m}\right]$, where $\alpha_{c}^{m}=e_{c}^{m}+1$. This formulation allows us to quantify the belief that the feature $v$ belongs to class $c$ as well as the uncertainty associated with the prediction as:
@@
The results from all prompt templates can be iteratively fused, as shown in the following equation:
$B_{c}^{m}$ =$=\begin{cases}b_{c}^{1},&\text{if }m=1\\ \frac{1}{1-C}(B_{c}^{m-1}b_{c}^{m}+B_{c}^{m-1}u^{m}+b_{c}^{m}U^{m-1}),&\text{if }1<m\leq M\end{cases},$ (8)
$U^{m}$ =$=\begin{cases}u^{1},&\text{if }m=1\\ \frac{1}{1-C}U^{m-1}u^{m},&\text{if }1<m\leq M\end{cases}.$
In this formulation, $B_{c}^{m}$ represents the belief after fusing the results from the first $m$ templates, while $U^{m}$ denotes the corresponding uncertainty. The process of evidence fusion can be understood as the summation of results over all semantics $s$ in Equation (2). Consequently, the result in Equation (2) can be reformulated as:
$P({\hat{Y}=c}|do({x}))=\frac{B_{c}^{M}}{\Sigma_{c^{\prime}=1}^{C}B_{c^{\prime}}^{M}}.$ (9)
To calculate the probability that sample $x$ belongs to class $c$ during the testing phase, we employ Equation (9). Accordingly, during the training stage, the cross-entropy loss which is used for learning the prompts is adjusted to the trusted cross-entropy loss to model the uncertainty information. For each training sample, the trusted cross-entropy can be formalized as:
@@ 4.3 Overall Architecture
where $\beta$ and $\gamma$ denote the hyper-parameters to tune the influence of $\mathcal{L}_{de}$ and $\mathcal{L}_{con}$, respectively. Refer to Algorithm 1 and Algorithm 2 for the detailed training and testing pipeline.
  repeat
     for $i$-$th$ training iteration do
        Iteratively sample a minibatch $\mathcal{X}^{\prime}$ from $\mathcal{X}$.
        for $m$-$th$ template do
           ${\mathcal{X}^{\prime}}^{m}=\mathcal{A}^{m}(\mathcal{X}^{\prime})$.
           Generate the text features $w^{m}$ with the text encoder.
           Generate the classification evidence $e^{m}$ with $v^{m}$ and $w^{m}$: $e_{c}^{m}=h(sim(w_{c}^{m},v^{m}))$.
           Using Equation (10) to calculate $\mathcal{L}_{t\text{-}ce}$.
           Using Equation (3) and Equation (5) to calculate $\mathcal{L}_{de}$ and $\mathcal{L}_{con}$.
           Using Equation (11) to calculate $\mathcal{L}_{CDC}$.
           $\theta^{m}=\theta^{m}-\ell\nabla_{\theta}\mathcal{L}_{CDC}$.
        end for
     end for
  until $\theta$ converge.
Algorithm 1 The training pipeline of CDC
  Sample a test data $x$ from $X$
  for $m$-$th$ template do
     Generate the image features $v^{m}$ of $x$.
     Generate the text features $w^{m}$ with the text encoder.
     Generate the classification evidence $e^{m}$ with $v^{m}$ and $w^{m}$: $e_{c}^{m}=h(sim(w_{c}^{m},v^{m}))$.
  end for
  Using Equation (8) and Equation (9) to generate the final classification results.
Algorithm 2 The testing pipeline of CDC
@@ 5 Experiments
Following previous works [5, 6], we conduct experiments to evaluate our proposed method with three different settings. These settings encompass the base-to-new setting as well as two out-of-distribution (OOD) settings, i.e., the cross-dataset setting and the cross-domain setting. Refer to Appendix B for a detailed overview of the evaluation protocol.
@@ 5.1 Experimental Settings
Experimental Details. We follow the experimental settings of the baseline method MaPLe. Specifically, we utilize a pre-trained CLIP with ViT-B/16 as the visual encoder. The number of learnable tokens is fixed at 2, whereas the prompt depth varies, being 9 for the base-to-new setting and 3 for the OOD setting. The learning rate is 0.035, and the batch size is 4. All models are trained using an SGD optimizer on an NVIDIA 3090 GPU. Our proposed CDC introduces three additional hyperparameters: $\beta$ and $\gamma$, which represent the weights for $\mathcal{L}_{de}$ and $\mathcal{L}_{con}$, respectively, and $M$, which denotes the number of prompts. Furthermore, different augmentation methods have a notable impact on the model’s performance. We set $\beta=5$, $\gamma=0.01$, and $M=4$ in the base-to-new setting, and $\beta=3$, $\gamma=0.01$, and $M=8$ in the OOD setting. For a detailed analysis of the influence of varying hyper-parameters and augmentation methods, please refer to Appendix C. We report the accuracies of base classes and novel classes, along with their harmonic mean (HM), averaged over 3 runs.